How do you solve the following problems using the substitution method? The end result should be a set of (x,y) coordinates.
2x+y=5
4x-3y=-10
4x+2y=7
x-5y=10
2x-y=2
x+2y=18.5
Thanks so much!

Answer by checkley77(12844)

2x+y=5 or y=5-2x
4x-3y=-10
4x-3(5-2x)=-10
4x-15+6x=-10
10x=-10+15
10x=5
x=5/10
x=.5 ans.
y=5-2*.5
y=5-1
y=4 ans.
--------------------------
4x+2y=7
x-5y=10 or x=5y+10
4(5y+10)+2y=7
20y+40+2y=7
22y=7-40
22y=33
y=33/22
y=3/2
y=1.5 ans.
x=5*1.5+10
y=7.5+10
y=17.5 ans
---------------------------
2x-y=2 or -y=-2x+2 or y=2x-2
x+2y=18.5
x+2(2x-2)=18.5
x+4x+4=18.5
5x=18.5-4
5x=14.5
x=14.5/5
x=2.9 ans.
y=2*2.9+2
y=5.8+2
y=7.8 ans.
-----------------------------